# How Much Do LLMs Hallucinate across Languages? On Multilingual Estimation of LLM Hallucination in the Wild

### v3

@@ Abstract
In the age of misinformation, hallucination—the tendency of Large Language Models (LLMs) to generate non-factual or unfaithful responses—represents the main risk for their global utility. Despite LLMs becoming increasingly multilingual, the vast majority of research on detecting and quantifying LLM hallucination are (a) English-centric and (b) focus on machine translation (MT) and summarization, tasks that are less common “in the wild” than open information seeking. In contrast, we aim to quantify the extent of LLM hallucination across languages in knowledge-intensive long-form question answering (LFQA). To this end, we train a multilingual hallucination detection model and conduct a large-scale study across 30 languages and 6 open-source LLM families. We start from an English hallucination detection dataset and rely on MT to translate-train a detection model. We also manually annotate gold data for five high-resource languages; we then demonstrate, for these languages, that the estimates of hallucination rates are similar between silver (LLM-generated) and gold test sets, validating the use of silver data for estimating hallucination rates for other languages. For the final rates estimation, we build open-domain QA dataset for 30 languages with LLM-generated prompts and Wikipedia articles as references. Our analysis shows that LLMs, in absolute terms, hallucinate more tokens in high-resource languages due to longer responses, but that the actual hallucination rates (i.e., normalized for length) seems uncorrelated with the sizes of languages’ digital footprints. We also find that smaller LLMs hallucinate more, and significantly, LLMs with broader language support display higher hallucination rates.
@@
On Multilingual Estimation of LLM Hallucination in the Wild
@@ 1 Introduction
Figure 1: Illustration of our approach for estimating hallucination rates in the wild. Hallucination Detection and Model Evaluation (left side): (1) We automatically translate the English FAVA Mishra et al. (2024) dataset to 30 languages and train our multilingual hallucination detection (HD) model on this (noisy) multilingual training data; (2) We synthesize a silver multilingual hallucination evaluation dataset by prompting a state-of-the-art LLM (GPT-4) to introduce hallucinations in its answers to knowledge-seeking questions; for a subset of five high-resource languages, we additionally collect gold (i.e., human) hallucination annotations; we dub this 30-language evaluation benchmark mFAVA. We use mFAVA to estimate HD model’s per-language performances (precision and recall). Hallucination Rate Estimation in the Wild (right side): (3) We estimate the hallucination rates for all 30 languages and six different LLM families from the number of detections of the HD model and its performance.
@@
Consequently, a large body of work on tackling LLM hallucination has emerged, with efforts falling into the three main areas: (1) detection, i.e., identification of the hallucinated content; (2) evaluation, primarily focusing on measures for quantifying the extent and severity of hallucinations ; and (3) mitigation, focusing on mitigating hallucinative tendencies of LLMs (Ji et al., 2023). While significant progress has been made in English (Maynez et al., 2020; Liu et al., 2022a; Obaid ul Islam et al., 2023; Kasai et al., 2024; Mishra et al., 2024), hallucination evaluation efforts targeting other languages have been much sparser Clark et al. (2023); Guerreiro et al. (2023); Herrlein et al. (2024); Shafayat et al. (2024). Moreover, these efforts have primarily targeted high-resource languages (Qiu et al., 2023; Shafayat et al., 2024) with benchmarks limited to reference-based tasks—text summarization (Clark et al., 2023; Aharoni et al., 2022) and machine translation (Dale et al., 2023; Guerreiro et al., 2023). While highly relevant, these tasks are arguably less representative of LLM usage ‘in the wild’ Trippas et al. (2024), where knowledge-intensive long-form question answering (LFQA) is more prominent.
@@
In this work, we address the above gaps in multilingual hallucination detection and evaluation research with the ultimate goal of estimating the “in the wild” hallucination rates of LLMs across languages. Multilingual estimation of such hallucination rates is challenging due to the scarcity of multilingual hallucination benchmarks covering open-ended knowledge-seeking tasks that are representative of real-world LLM usage: unlike in reference-based generation tasks like summarization and machine translation, LLMs often generate long-form responses to open-ended questions, requiring more comprehensive evaluation approaches (Wei et al., 2024b). Concretely, we present a large-scale study that estimates hallucination rates for 30 languages (both high(er)- and low(er)-resource languages). Our main contributions are as follows: (1) We translate-train Artetxe et al. (2023); Ebing and Glavaš (2024) a multilingual hallucination detection (HD) model on 30 languages. (2) We create mFAVA HD evaluation datasets with span-level human annotations (mFAVA-gold) for five high-resource languages, generate synthetic (mFAVA-silver) HD evaluation datasets for 25 additional languages, and validate the use of mFAVA-silver by showing the mFAVA-silver and mFAVA-gold estimates yield similar results; (3) We propose a protocol for estimating “in the wild” hallucination rates of LLMs and introduce an extensive synthetic dataset (51,133 prompts across 30 languages) for estimating LLM hallucination rates in highly multilingual settings; (4) We offer a comprehensive hallucination rate analysis of six LLM families, validating previous findings that larger models tend to hallucinate less, and uncovering that broader LLM language coverage correlates with increased hallucination rates. This work is the first to estimate “in-the-wild” LLM hallucination rates for a wide range of languages using knowledge-intensive open-domain LFQA, reflecting real-world usage. Our comprehensive framework is illustrated in Figure 1111We release our datasets and work on: https://github.com/WorldHellow/mHallucinations-LLM.
@@ 3 Hallucination Detection
Binary Multi Bidirect 89.5 65.0 69.7 58.7 82.5 61.6 89.1 65.5 86.4 72.5
Multi Causal 81.8 59.6 76.3 62.2 75.3 60.0 75.8 55.6 75.7 67.3
@@
Category Multi Bidirect 73.2 45.0 46.5 30.1 66.1 37.2 72.3 41.5 72.9 51.8
Multi Causal 68.7 43.4 56.5 34.1 51.8 29.4 62.6 37.9 58.6 42.4
Table 3: Token-level F1 performance of multilingual (Multi) and monolingual (Mono) hallucination detection models for five high-resource languages with both Silver and Gold evaluation data in mFAVA. Performance reported for hallucination detection alone (Binary) and hallucination detection and type classification (Category). Models fine-tuned without (Bidirect) or with (Causal) future token masking. Bold: best result in each column.
@@ 3.2 Multilingual Hallucination Detection
Results. Table 3 summarizes the hallucination detection performance for five high-resource languages for which we have both LLM-synthesized Silver data and human-annotated Gold portions in our mFAVA benchmark. We first observe that, expectedly, just detecting hallucinated spans (Binary task) is much easier than additionally correctly recognizing the type of hallucination (Category task). Although category labels offer finer-grained insight into the nature of LLM hallucination, we deem the models’ performance on fine-grained hallucination type classification—especially on Gold, human-labeled portions of mFAVA—insufficient for reliably estimating type-specific hallucination rates “in the wild” (see §4.2). These results are in line with IAA from Figure 2, with consistently larger IAA for hallucination detection (Binary) then for type classification (Category). This renders fine-grained hallucination type classification difficult for both humans and models and warrants a broader research effort on hallucination type taxonomies as well as better hallucination type detection models. We leave this for future work.
Models’ performance on the detection-only (Binary) tasks is much better across the board, but the results are much better on the Silver portions (hallucinations generated by GPT-4) of mFAVA than on the Gold (human-labeled hallucination spans). This is expected, because the hallucinated spans in our training data have also been generated by GPT-4—this means that the human-annotated Gold mFAVA portions introduce much more of a distribution shift w.r.t. training data than the corresponding Silver portions. At this point it is important to (re-)emphasize that we are not really interested in the absolute performance of the detection models, but rather using these detection performance estimates to produce reliable hallucination rate estimates for LLMs in the wild (§4.2).
We next observe that the 30-language multilingual model (Multi) is consistently better than language-specific monolingual models (Mono), with gaps being particularly wide in the Category task (e.g., +30 F1 points for Arabic on the Gold mFAVA portion). Albeit smaller, the differences are also substantial in the Binary hallucination detection (e.g., +7 F1 points for Arabic and German, on respective Gold mFAVA portions). Finally, bidirectional contextualization in fine-tuning (Bidirect) seems to be generally more effective than fine-tuning with future-token masking (Causal), with Chinese performance as the only exception. This is in line with findings from other token-classification tasks
@@
4 Estimating Hallucination in the Wild
We next propose a protocol for estimating hallucination rates of LLMs (for a wide range of languages) in the wild, based (1) on the number of hallucinated tokens detected by a hallucination detection (HD) model in the wild and (2) estimates of HD model’s performance (precision and recall).
@@ 4.1 From Model Performance to Hallucination Rates Estimates
Estimating Hallucination Rates in the Wild. Let $\mathit{P}_{l}$ and $\mathit{R}_{l}$ be the estimates of token-level precision and recall of a HD model for some language $l$ and let $\mathit{H}_{\textit{det},l}$ be the number of hallucination tokens that the HD model detected (i.e., predicted) on some corpus $C_{l}$ of LLM generations in language $l$, which serves as an approximation of the LLM outputs in the wild. We then posit that the estimate of the true hallucination rate of the LLM in the wild for language $l$, $\mathit{HR}_{\text{est},l}$, is given as follows:
@@
Estimation Dataset. We next create corpora $C_{l}$ (one corpus for each of our 30 target languages) of free-text LLM answers to knowledge-intensive queries, as approximations of the LLM usage in the wild. We start by randomly selecting articles from the language-specific Wikipedia, to serve as ground truth reference text. To ensure quality of reference text, we choose only from Wikipedia articles that are at least 2,000 characters long and have the collaborative Wikipedia depth (Alshahrani et al., 2023) of at least 5.555The depth indicates the number of collaborative edits and correlates with the quality/factuality of the content. We then prompt GPT-4 to generate two knowledge-intensive queries for each selected article, ensuring that the information required to answer to the query is fully contained in the article text (see Table 11 in the §A.5 for the exact prompt). As a sanity check, we manually checked for 50 synthesized queries and five languages from Table 3—by translating the query and reference article to English—whether the answers to queries are indeed contained in the article, establishing that this is indeed so in 98% of cases. Our final dataset for multilingual hallucination rate estimation consists of 25,685 Wikipedia articles (spanning over 15,940 unique Wikipedia categories) and 51,133 queries. Table 9 in §A provides per-language statistics. We provide details on constructing the datasets in §A.5.
@@
Estimates from mFAVA-Silver Performance. On the one hand, creating Gold datasets for hallucination detection evaluation is prohibitively expensive (see §A.2)—this is why we obtained such annotations for only five of 30 mFAVA languages. On the other hand, the estimates of HD model’s performance are much higher on mFAVA-Silver (see Table 3), with GPT-4-labeled hallucinations: this, at first glance, questions the validity of estimating ‘in the wild’ hallucination rates based on $\mathit{P}_{l}$ and $\mathit{R}_{l}$ estimated on Silver data, for the 25 languages for which we do not have mFAVA-Gold portions. Recall, however, that we do not care about HD model’s absolute $\mathit{P}_{l}$ and $\mathit{R}_{l}$, but whether the $\mathit{P}_{l}$ and $\mathit{R}_{l}$ estimates can produce reliable hallucination rate estimates $\mathit{HR}_{\text{est},l}$. Looking at Eq. 1, $\mathit{HR}_{\text{est},l}$ depends on the ratio $\mathit{P}_{l}/\mathit{R}_{l}$ and not absolute values of $\mathit{P}_{l}$ and $\mathit{R}_{l}$. We thus next test, for the five languages with both Silver and Gold portions in mFAVA, whether the $\mathit{HR}_{\text{est},l}$ estimates based on the Silver $\mathit{P}_{l}$ and $\mathit{R}_{l}$ (roughly) match those based on Gold $\mathit{P}_{l}$ and $\mathit{R}_{l}$. Figure 3 shows $\mathit{HR}_{\text{est},l}$ estimates, computed from the performance of our Multi (Bidirect) model on Silver and Gold portions, respectively, and number of its hallucination detections $\mathit{H}_{\textit{det},l}$ on outputs of three LLMs: Llama-3-8B, Qwen-2.5-7B, and Aya-8B. We observe very strong Pearson correlation (r=0.83,p=1.26e04)(r=0.83,p=1.26e-04) between the Gold-based and Silver-based $\mathit{HR}_{\text{est},l}$ estimates, which, we argue, justifies the usage of Silver mFAVA datasets for estimating $\mathit{HR}_{\text{est},l}$ for the 25 languages without the Gold mFAVA portions.

### v4

# How Much Do LLMs Hallucinate across Languages? On Realistic Multilingual Estimation of LLM Hallucination

@@ Abstract
In the age of misinformation, hallucination—the tendency of Large Language Models (LLMs) to generate non-factual or unfaithful responses—represents the main risk for their global utility. Despite LLMs becoming increasingly multilingual, the vast majority of research on detecting and quantifying LLM hallucination are (a) English-centric and (b) focus on machine translation (MT) and summarization, tasks that are less common in realistic settings than open information seeking. In contrast, we aim to quantify the extent of LLM hallucination across languages in knowledge-intensive long-form question answering (LFQA). To this end, we train a multilingual hallucination detection model and conduct a large-scale study across 30 languages and 6 open-source LLM families. We start from an English hallucination detection dataset and rely on MT to translate-train a detection model. We also manually annotate gold data for five high-resource languages; we then demonstrate, for these languages, that the estimates of hallucination rates are similar between silver (LLM-generated) and gold test sets, validating the use of silver data for estimating hallucination rates for other languages. For the final rates estimation, we build open-domain QA dataset for 30 languages with LLM-generated prompts and Wikipedia articles as references. Our analysis shows that LLMs, in absolute terms, hallucinate more tokens in high-resource languages due to longer responses, but that the actual hallucination rates (i.e., normalized for length) seems uncorrelated with the sizes of languages’ digital footprints. We also find that smaller LLMs hallucinate more, and significantly, LLMs with broader language support display higher hallucination rates.
@@
On Realistic Multilingual Estimation of LLM Hallucination
@@
Figure 1: Illustration of our approach for estimating hallucination rates. Hallucination Detection and Model Evaluation (left side): (1) We automatically translate the English FAVA Mishra et al. (2024) dataset to 30 languages and train our multilingual hallucination detection (HD) model on this (noisy) multilingual training data; (2) We synthesize a silver multilingual hallucination evaluation dataset by prompting a state-of-the-art LLM (GPT-4) to introduce hallucinations in its answers to knowledge-seeking questions; for a subset of five high-resource languages, we additionally collect gold (i.e., human) hallucination annotations; we dub this 30-language evaluation benchmark mFAVA. We use mFAVA to estimate HD model’s per-language performances (precision and recall). Hallucination Rate Estimation (right side): (3) We estimate the hallucination rates for all 30 languages and six different LLM families from the number of detections of the HD model and its performance.
@@ 1 Introduction
Consequently, a large body of work on tackling LLM hallucination has emerged, with efforts falling into the three main areas: (1) detection, i.e., identification of the hallucinated content; (2) evaluation, primarily focusing on measures for quantifying the extent and severity of hallucinations ; and (3) mitigation, focusing on mitigating hallucinative tendencies of LLMs (Ji et al., 2023). While significant progress has been made in English (Maynez et al., 2020; Liu et al., 2022a; Obaid ul Islam et al., 2023; Kasai et al., 2024; Mishra et al., 2024), hallucination evaluation efforts targeting other languages have been much sparser Clark et al. (2023); Guerreiro et al. (2023); Herrlein et al. (2024); Shafayat et al. (2024). Moreover, these efforts have primarily targeted high-resource languages (Qiu et al., 2023; Shafayat et al., 2024) with benchmarks limited to reference-based tasks—text summarization (Clark et al., 2023; Aharoni et al., 2022) and machine translation (Dale et al., 2023; Guerreiro et al., 2023). While highly relevant, these tasks are arguably less representative of LLM usage in realistic scenarios Trippas et al. (2024), where knowledge-intensive long-form question answering (LFQA) is more prominent.
@@
In this work, we address the above gaps in multilingual hallucination detection and evaluation research with the ultimate goal of estimating the hallucination rates of LLMs across languages in a realistic scenario of open-domain QA. Multilingual estimation of such hallucination rates is challenging due to the scarcity of multilingual hallucination benchmarks covering open-ended knowledge-seeking tasks that are representative of real-world LLM usage: unlike in reference-based generation tasks like summarization and machine translation, LLMs often generate long-form responses to open-ended questions, requiring more comprehensive evaluation approaches (Wei et al., 2024b). Concretely, we present a large-scale study that estimates hallucination rates for 30 languages (both high(er)- and low(er)-resource languages). Our main contributions are as follows: (1) We translate-train Artetxe et al. (2023); Ebing and Glavaš (2024) a multilingual hallucination detection (HD) model on 30 languages. (2) We create mFAVA HD evaluation datasets with span-level human annotations (mFAVA-gold) for five high-resource languages, generate synthetic (mFAVA-silver) HD evaluation datasets for 25 additional languages, and validate the use of mFAVA-silver by showing the mFAVA-silver and mFAVA-gold estimates yield similar results; (3) We propose a protocol for estimating hallucination rates in open domain LFQA of LLMs and introduce an extensive synthetic dataset (51,133 prompts across 30 languages) for estimating LLM hallucination rates in highly multilingual settings; (4) We offer a comprehensive hallucination rate analysis of six LLM families, validating previous findings that larger models tend to hallucinate less, and uncovering that broader LLM language coverage correlates with increased hallucination rates. This work is the first to estimate LLM hallucination rates for a wide range of languages using knowledge-intensive open-domain LFQA, reflecting real-world usage. Our comprehensive framework is illustrated in Figure 1111We release our datasets and work on: https://github.com/WorldHellow/mHallucinations-LLM.
@@ 3 Hallucination Detection
Binary Multi Bidirect 89.5* 65.0* 69.7 58.7 82.5* 61.6 89.1* 65.5* 86.4* 72.5*
Multi Causal 81.8 59.6 76.3* 62.2* 75.3 60.0 75.8 55.6 75.7 67.3
@@
Category Multi Bidirect 73.2* 45.0 46.5 30.1 66.1* 37.2* 72.3* 41.5* 72.9* 51.8*
Multi Causal 68.7 43.4 56.5* 34.1* 51.8 29.4 62.6 37.9 58.6 42.4
Table 3: Token-level F1 performance of multilingual (Multi) and monolingual (Mono) hallucination detection models for five high-resource languages with both Silver and Gold evaluation data in mFAVA. Performance reported for hallucination detection alone (Binary) and hallucination detection and type classification (Category). Models fine-tuned without (Bidirect) or with (Causal) future token masking. Bold: best result in each column, Asterisk: significantly higher ($p<0.05$) score.
@@ 3.2 Multilingual Hallucination Detection
Results. Table 3 summarizes the hallucination detection performance for five high-resource languages for which we have both LLM-synthesized Silver data and human-annotated Gold portions in our mFAVA benchmark. We first observe that, expectedly, just detecting hallucinated spans (Binary task) is much easier than additionally correctly recognizing the type of hallucination (Category task). Although category labels offer finer-grained insight into the nature of LLM hallucination, we deem the models’ performance on fine-grained hallucination type classification—especially on Gold, human-labeled portions of mFAVA—insufficient for reliably estimating type-specific hallucination rates (see §4.2). These results are in line with IAA from Figure 2, with consistently larger IAA for hallucination detection (Binary) then for type classification (Category). This renders fine-grained hallucination type classification difficult for both humans and models and warrants a broader research effort on hallucination type taxonomies as well as better hallucination type detection models. We leave this for future work.
Models’ performance on the detection-only (Binary) tasks is much better across the board, but the results are much better on the Silver portions (hallucinations generated by GPT-4) of mFAVA than on the Gold (human-labeled hallucination spans). This is expected, because the hallucinated spans in our training data have also been generated by GPT-4—this means that the human-annotated Gold mFAVA portions introduce much more of a distribution shift w.r.t. training data than the corresponding Silver portions. At this point it is important to (re-)emphasize that we are not really interested in the absolute performance of the detection models, but rather using these detection performance estimates to produce reliable hallucination rate estimates for LLMs in realistic setting (§4.2).
We next observe that the 30-language multilingual model (Multi) is consistently better than language-specific monolingual models (Mono), with gaps being particularly wide in the Category task (e.g., +30 F1 points for Arabic on the Gold mFAVA portion). Albeit smaller, the differences are also substantial in the Binary hallucination detection (e.g., +7 F1 points for Arabic and German, on respective Gold mFAVA portions). Finally, bidirectional contextualization in fine-tuning (Bidirect) seems to be generally more effective than fine-tuning with future-token masking (Causal), with Chinese performance as the only exception. This is in line with findings from other token-classification tasks Li et al. (2023b); Dukić and Šnajder (2024).
@@
4 Estimating Hallucination in Realistic Setting
We next propose a protocol for estimating hallucination rates of LLMs (for a wide range of languages), based (1) on the number of hallucinated tokens detected by a hallucination detection (HD) model in the wild and (2) estimates of HD model’s performance (precision and recall).
@@ 4.1 From Model Performance to Hallucination Rates Estimates
Estimating Hallucination Rates. Let $\mathit{P}_{l}$ and $\mathit{R}_{l}$ be the estimates of token-level precision and recall of a HD model for some language $l$ and let $\mathit{H}_{\textit{det},l}$ be the number of hallucination tokens that the HD model detected (i.e., predicted) on some corpus $C_{l}$ of LLM generations in language $l$, which serves as an approximation of the LLM outputs in the wild. We then posit that the estimate of the true hallucination rate of the LLM in the wild for language $l$, $\mathit{HR}_{\text{est},l}$, is given as follows:
@@
Estimation Dataset. We next create corpora $C_{l}$ (one corpus for each of our 30 target languages) of free-text LLM answers to knowledge-intensive queries, as approximations of the LLM usage in the real world. We start by randomly selecting articles from the language-specific Wikipedia, to serve as ground truth reference text. To ensure quality of reference text, we choose only from Wikipedia articles that are at least 2,000 characters long and have the collaborative Wikipedia depth (Alshahrani et al., 2023) of at least 5.555The depth indicates the number of collaborative edits and correlates with the quality/factuality of the content. We then prompt GPT-4 to generate two knowledge-intensive queries for each selected article, ensuring that the information required to answer to the query is fully contained in the article text (see Table 11 in the §A.5 for the exact prompt). As a sanity check, we manually checked for 50 synthesized queries and five languages from Table 3—by translating the query and reference article to English—whether the answers to queries are indeed contained in the article, establishing that this is indeed so in 98% of cases. Our final dataset for multilingual hallucination rate estimation consists of 25,685 Wikipedia articles (spanning over 15,940 unique Wikipedia categories) and 51,133 queries. Table 9 in §A provides per-language statistics. We provide details on constructing the datasets in §A.5.
@@
Estimates from mFAVA-Silver Performance. On the one hand, creating Gold datasets for hallucination detection evaluation is prohibitively expensive (see §A.2)—this is why we obtained such annotations for only five of 30 mFAVA languages. On the other hand, the estimates of HD model’s performance are much higher on mFAVA-Silver (see Table 3), with GPT-4-labeled hallucinations: this, at first glance, questions the validity of estimating realistic hallucination rates based on $\mathit{P}_{l}$ and $\mathit{R}_{l}$ estimated on Silver data, for the 25 languages for which we do not have mFAVA-Gold portions. Recall, however, that we do not care about HD model’s absolute $\mathit{P}_{l}$ and $\mathit{R}_{l}$, but whether the $\mathit{P}_{l}$ and $\mathit{R}_{l}$ estimates can produce reliable hallucination rate estimates $\mathit{HR}_{\text{est},l}$. Looking at Eq. 1, $\mathit{HR}_{\text{est},l}$ depends on the ratio $\mathit{P}_{l}/\mathit{R}_{l}$ and not absolute values of $\mathit{P}_{l}$ and $\mathit{R}_{l}$. We thus next test, for the five languages with both Silver and Gold portions in mFAVA, whether the $\mathit{HR}_{\text{est},l}$ estimates based on the Silver $\mathit{P}_{l}$ and $\mathit{R}_{l}$ (roughly) match those based on Gold $\mathit{P}_{l}$ and $\mathit{R}_{l}$. Figure 3 shows $\mathit{HR}_{\text{est},l}$ estimates, computed from the performance of our Multi (Bidirect) model on Silver and Gold portions, respectively, and number of its hallucination detections $\mathit{H}_{\textit{det},l}$ on outputs of three LLMs: Llama-3-8B, Qwen-2.5-7B, and Aya-8B. We observe very strong Pearson correlation (r=0.83,p=1.26e04)(r=0.83,p=1.26e-04) between the Gold-based and Silver-based $\mathit{HR}_{\text{est},l}$ estimates, which, we argue, justifies the usage of Silver mFAVA datasets for estimating $\mathit{HR}_{\text{est},l}$ for the 25 languages without the Gold mFAVA portions.
@@ 4.2 Final Estimates
Figure 6: Significant interaction effects from the linear mixed effects model. (a) Model Size $\times$ Language Support interaction ($\beta=1.33$, $p<0.001$). (b) Model Size $\times$ Response Length interaction ($\beta=-1.02$, $p=0.001$). Dots represent observed means; shaded areas show the magnitude of difference between model sizes.
4.3 Mixed Effect Modeling
To understand the potential dependencies between factors identified in §4.2, we conducted a linear mixed effects analysis with hallucination rate as the dependent variable. The model included fixed effects for model size (small: 1.7-3B vs. large: 7-9B parameters), number of supported languages, and response length, along with all two-way interactions. The interaction model significantly improved fit over the main effects model ($\text{Likelihood Ratio}=22.14$, $p<0.001$), showing that these factors do not operate independently.
Figure 6 reveals two significant interactions: (1) Model Size $\times$ Language Support ($\beta=1.33$, $p<0.001$): the hallucination gap between small and large models widens as language support increases—small models struggle disproportionately with multilingual capabilities; and (2) Model Size $\times$ Response Length ($\beta=-1.02$, $p=0.001$): larger models excel particularly for longer responses, while this advantage diminishes for shorter outputs. These findings reveal that model size acts as a critical moderator—larger models not only hallucinate less overall but are also more robust to the challenges posed by multilingual support and longer generation tasks, suggesting that scaling model parameters provides compound benefits beyond simple performance improvements.